OpenStudy (bbb911):

x - 7/x2 - 49 Find all numbers for which the rational expression is not defined.

OpenStudy (anonymous):

hey :)

OpenStudy (goformit100):

▬Sir/Ma'am, A Warm Welcome To Open Study. Sir/Ma'am please show your Working for the question you have posted.▬

OpenStudy (bbb911):

hi :)

OpenStudy (bbb911):

spamming again

OpenStudy (anonymous):

@goformit100 dont worry :) ill help her

OpenStudy (bbb911):

:)

OpenStudy (anonymous):

ight lets take a look at the question :)

OpenStudy (bbb911):

okay

OpenStudy (anonymous):

ok so all you want is me to solve the equation?

OpenStudy (bbb911):

help me out with each step im not getting it

OpenStudy (anonymous):

lol ok no prob to help and explain the steps right?

OpenStudy (bbb911):

yes thank you :)

OpenStudy (anonymous):

ok :)

OpenStudy (elsa123):

Final result : x3 - 49x2 + 7 ————————————— x2 Step by step solution : Step 1 : 7 Simplify x + —— x2 Rewriting the whole as an Equivalent Fraction : 1.1 Adding a fraction to a whole Rewrite the whole as a fraction using x2 as the denominator : x x • x2 x = — = —————— 1 x2 Equivalent fraction : The fraction thus generated looks different but has the same value as the whole Common denominator : The equivalent fraction and the other fraction involved in the calculation share the same denominator Adding fractions that have a common denominator : 1.2 Adding up the two equivalent fractions Add the two equivalent fractions which now have a common denominator Combine the numerators together, put the sum or difference over the common denominator then reduce to lowest terms if possible: x • x2 + 7 x3 + 7 —————————— = —————— x2 x2 Equation at the end of step 1 : (x3 + 7) ———————— - 49 x2 Step 2 : x3+7 Simplify ———— - 49 x2 Rewriting the whole as an Equivalent Fraction : 2.1 Subtracting a whole from a fraction Rewrite the whole as a fraction using x2 as the denominator : 49 49 • x2 49 = —— = ——————— 1 x2 Trying to factor as a Sum of Cubes : 2.2 Factoring: x3 + 7 Theory : A sum of two perfect cubes, a3 + b3 can be factored into : (a+b) • (a2-ab+b2) Proof : (a+b) • (a2-ab+b2) = a3-a2b+ab2+ba2-b2a+b3 = a3+(a2b-ba2)+(ab2-b2a)+b3= a3+0+0+b3= a3+b3 Check : 7 is not a cube !! Ruling : Binomial can not be factored as the difference of two perfect cubes Polynomial Roots Calculator : 2.3 Find roots (zeroes) of F(x) = x3 + 7 Polynomial Roots Calculator is a set of methods aimed at finding values of x for which F(x)=0 Rational Roots Test is one of the above mentioned tools. It would only find Rational Roots that is numbers x which can be expressed as the quotient of two integers The Rational Root Theorem states that if a polynomial zeroes for a rational number P/Q then P is a factor of the Trailing Constant and Q is a factor of the Leading Coefficient In this case, the Leading Coefficient is 1 and the Trailing Constant is 7. The factor(s) are: of the Leading Coefficient : 1 of the Trailing Constant : 1 ,7

OpenStudy (bbb911):

thank u

OpenStudy (elsa123):

may i have medal :)

OpenStudy (bbb911):

yes i am going to repost this and just comment on it like ur welcome and ill give you a medal

OpenStudy (anonymous):

its cool bbb91 :) you can give it to her :)